A mixture of 0.4500 M CO and 0.4450 M Cl2 is enclosed in a vessel and heated to 1000K .

CO(g)+Cl2(g)↽−−⇀COCl2(g) ?c=255.0 at 1000

Calculate the equilibrium concentration of each gas at 1000 K1000 K .
Answer #1

                        CO(g) + Cl2(g)--------->COCl2(g)

        I               0.45       0.445                 0

        C               -x              -x                   +x

         E            0.45-x       0.445-x               +x

                        Kc     =   [COCl2]/[CO][Cl2]

                         255     = x/(0.45-x)(0.445-x)

                        255*(0.45-x)(0.445-x) = x

                          x   = 0.407

           [CO]    = 0.45-x   = 0.45-0.407   = 0.043M

           [Cl2]     = 0.445-x   = 0.445-0.407   = 0.038M

          [COCl2]    = x          = 0.407M